A local hamburger shop sold a combined total of 697 hamburgers and cheeseburgers on Thursday. There were 53 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Thursday?
Answer by oberobic(2304)

h = hamburgers
c = cheeseburgers
h + c = 697
so
h = 697-c
.
h = c+53
.
h=h
so
697-c = c+53
-2c = -644
c = 322
.
h = 697 - 322
h = 375
.
Answer: 375 hamburgers and 322 cheeseburgers
.
Done.